# Partonomy: Large Multimodal Models with Part-Level Visual Understanding

### v2

@@ Abstract
Real-world objects are composed of distinctive, object-specific parts. Identifying these parts is key to performing fine-grained, compositional reasoning—yet, large multimodal models (LMMs) struggle to perform this seemingly straightforward task. In this work, we introduce Partonomy, an LMM benchmark designed for pixel-level part grounding. We construct Partonomy from existing part datasets and our own rigorously annotated set of images, encompassing 862 part labels and 534 object labels for evaluation. Unlike existing datasets that simply ask models to identify generic parts, Partonomy uses specialized concepts (e.g., agricultural airplane), and challenges models to compare objects’ parts, consider part-whole relationships, and justify textual predictions with visual segmentations. Our experiments demonstrate significant limitations in state-of-the-art LMMs (e.g., LISA-13B achieves only 5.9% gIoU), highlighting a critical gap in their part grounding abilities. We note that existing segmentation-enabled LMMs (segmenting LMMs) have two key architectural shortcomings: they use special [SEG] tokens not seen during pretraining which induce distribution shift, and they discard predicted segmentations instead of using past predictions to guide future ones. To address these deficiencies, we train several part-centric LMMs and propose Plum, a novel segmenting LMM that uses span tagging instead of segmentation tokens and that conditions on prior predictions in a feedback loop. We find that pretrained Plum outperforms existing segmenting LMMs on reasoning segmentation, VQA, and visual hallucination benchmarks. In addition, Plum finetuned on our proposed Explanatory Part Segmentation task is competitive with segmenting LMMs trained on significantly more segmentation data. Our work opens up new avenues towards enabling fine-grained, grounded visual understanding in LMMs. The code and data will be made publicly available at: https://github.com/AnselBlume/partonomy
@@ 1 Introduction
Figure 1: Example of the Explanatory Part Segmentation task in Partonomy and its construction. Given an input image, segmentation-enabled LMMs should generate text explanations and part-based segmentation masks, which respectively serve as textual and visual rationales for their decision. Our question-answer mutator generates challenging answer choices based on the part set overlap between object-level concepts.
Real-world objects can be decomposed into distinctive parts. A banana boat (Fig. 2), for instance, consists of a seating tube, a handle, a hull, and an inflation valve. Such distinctive parts characterize each concept, differentiating one object from another. The ability to recognize and distinguish between parts is an important element of holistic object understanding, with applications ranging from explainable object recognition blume2024miracle ; kim-ji-2024-finer ; liu2023llava ; dai2023instructblip ; zhu2023minigpt , to part-based novel concept design design2025 , and robotic manipulation xu2023creativerobottooluse ; liu2024composable ; jiang2024robot . Decomposing objects into their key building blocks allows models to reason about objects at a granular level kim-ji-2024-finer ; liu2024democratizing , allowing for more complex and nuanced interactions.
@@
To quantify the part recognition abilities of LMMs, we propose Explanatory Part Segmentation, a task that assesses LMMs’ ability to recognize object parts, associate objects with their distinctive parts, and use these grounded parts to predict object labels. We then introduce Partonomy, a comprehensive benchmark for the Explanatory Part Segmentation task. We construct Partonomy from existing part segmentation datasets he2022partimagenet ; chen2014pascalpart ; ramanathan2023paco and our manually-annotated evaluation dataset of 1K specialized object-centric images with complex part annotations. This subset, Partonomy-Core, contains 862 distinct part labels—more than any existing part datasets (Section 3.1).
We then note two shortcomings of existing segmenting LMMs’ architectures lai2024lisa ; rasheed2024glamm ; ren2024pixellm . First, they always rely on special [SEG] tokens not seen during pretraining, potentially hindering downstream performance by introducing distribution shift. Second, these segmenting LMMs discard their predictions after each output, missing the opportunity to incorporate prior information contained by the masks they predicted during the decoding process. This design is in contrast to modern generative frameworks, which condition future predictions on past ones vaswani2017attention ; sohl2015thermodynamics . Based on these observations, we propose Plum, a Part-Level Understanding LMM. Plum uses a text span tagging module to avoid special segmentation tokens that induce distribution shift from the pre-trained LLM, and employs a mask feedback mechanism to condition on past predictions (Section 4). Our results show that pretrained Plum retains its general reasoning abilities far better than other segmenting LMMs, achieving stronger zero-shot segmentation performance and competitive finetuned performance to models trained on significantly more segmentation data.
@@ 2 Related Work
Figure 2: An example of pretrained PLUM’s part understanding compared to recent segmenting LMMs trained on part data. The other models do not segment any part of the image.
LMMs such as LISA (lai2024lisa, ) and GLaMM (rasheed2024glamm, ) have demonstrated the ability to generate text and grounded segmentation masks. Despite being trained on part-level segmentation datasets such as PACO (ramanathan2023paco, ) and Pascal-Part (pascal-voc-2007, ), they struggle to exhibit a part-level understanding of visual concepts. While they demonstrate the ability to understand complex textual instructions (lai2024lisa, ; rasheed2024glamm, ; ren2024pixellm, ; yuan2024osprey, ), current LMMs fail to relate concept-indicative parts (e.g., seating tube) to their wholes (e.g., banana boat), as shown in Fig. 2; particularly, they fail to generate the specialized token (e.g., [SEG]) added to their vocabulary, leading to no masks being generated for the parts. Even state-of-the-art LMMs that seemingly “reason” struggle to establish attributive relationships between objects and parts, implying that current models and datasets lack the coverage and capacity to handle complex part understanding and grounding tasks. This observation motivates the proposal of our new Explanatory Part Segmentation task, which requires LMMs to segment objects’ parts (i.e. producing visual rationale) while generating the corresponding text rationale.
@@ 3.1 Task Overview
To evaluate these elements of part understanding, we define the Explanatory Part Segmentation task. In this task, a model is provided with an image and a question about an object’s parts (e.g. “What visible parts does the agricultural plane in the image have?” Fig. 1). The model must then select the best response and generate segmentation masks for the corresponding parts to explain its selection (e.g. “The agricultural plane has wings, a propulsion component, and a spraying rig”). Motivated by our characterization of part understanding, we define three classes of questions (Fig. 1):
@@
{tblr}
colspec = p2.9cm c c c c c c c, row1 = font=, width = Datasets # Object Labels # Part Labels # Object-Part Labels # Images # Seg. Masks Object-Centric Images Object Domain
PASCAL-Part 20 30 193 10,103 111,960 ✗ C, O
PartImageNet 158 14 14 24,000 112,000 ✓ C, O
PartImageNet++ 1000 818 3,308 100K 406.4K ✓ C, O
PACO 75 200 456 84,027 641,000 ✗ C
Partonomy -Core 534 862 1,976 1,068 4,968 ✓ C, O, S
@@ 3.3 Evaluation
We evaluate answer selection via accuracy (random $=$ 20% for 5 answer choices) for both part prediction (all questions) and object prediction (Part-Whole Reasoning questions). However, some wrong answer choices are better than others. Precision and recall capture the similarity of the predicted parts to the ground truth set, and we adopt these as more fine-grained measurements of part recognition by the LMMs.
@@ 4 Plum: Part-Level Understanding LMM
We find that existing LMMs are unable to accurately identify parts in an image. Even segmenting LMMs (lai2024lisa, ; rasheed2024glamm, ; ren2024pixellm, ) trained on part segmentation datasets such as PACO (ramanathan2023paco, ) and Pascal-Part (chen2014pascalpart, ) exhibit poor performance on part-level segmentation (Figs. 4 and 2). We identify two key architectural deficiencies of segmenting LMMs: (1) They rely on special tokens for segmentation (e.g. [SEG] or <p></p>). These tokens are not seen during pretraining, so we hypothesize that their addition to the vocabulary and subsequent finetuning perturbs models’ original token distributions (Table 5). (2) They discard prior mask predictions when segmenting in sequence, conditioning only on past text during generation. Incorporating prior mask predictions would likely help maintain consistency and better localize future predictions (Fig. 5(a)).
@@
Let $h_{i}^{L}\!\in\!\mathbb{R}^{d}$ be the VLM’s last-layer embedding of token $i$ $(i=1,\dots,N)$ of the output sequence. We process these embeddings along two complementary pathways: (i) the Span Extractor, a bidirectional self-attention block that tags beginning (B), inside (I), and outside (O) (ramshaw1999text, ) positions of tokens to segment, and (ii) a projection head that maps B/I embeddings into “mask queries,” regularized by KL divergence. An overview is given in Fig. 3.
@@
[…]
@@
For each B or I span token, we project $q_{i}=g(h_{i}^{L})$ and impose a Gaussian KL penalty $\small\mathcal{L}_{\text{KL}}=\frac{1}{N_{+}}\sum_{y_{i}\neq\textsc{O}}\frac{\lVert h_{i}^{L}-q_{i}\rVert^{2}}{2\sigma^{2}}$ which is applied only to B/I tokens to prevent their hidden states from drifting away from the pre-trained text representation space. As we see later, this is crucial to retain the textual knowledge and reasoning ability of the pretrained LMM.
@@
To incorporate previously predicted masks into the mask decoding process, we inject feature-wise linear modulation (FiLM) (perez2018film, ) layers into the SAM decoder’s mask encoder (Fig. 3). These layers allow us to encode the mask while conditioning on prior text spans, providing semantics beyond a raw binary mask. We use this modified mask encoder to encode each prior mask into a stack of text-enhanced feature maps. The stack of feature maps is pooled into a single feature map via patch-wise attention pooling (over the stack dimension), with a learned feature map providing an attentional query for each patch. This pooled feature map representing all prior predicted masks is fed into the mask decoder (along with the pooled text embeddings) to generate the next mask.
@@
Tagged token embeddings are average pooled and passed to the mask decoder, generating a mask $\hat{M}_{i}$. With ground-truth mask $M_{i}$ we adopt the Focal-Tversky loss (abraham2019novel, ), $\mathcal{L}_{\text{seg}}=\frac{1}{N_{+}}\!\sum_{y_{i}\neq\textsc{O}}\mathcal{L}_{\text{FT}}(M_{i},\hat{M}_{i})$. Focal-Tversky loss is a generalized version of the DICE loss (sudre2017generalised, ). This gives the overall objective equal to $\mathcal{L}=\mathcal{L}_{\text{LM}}+\lambda_{\text{1}}\mathcal{L}_{\text{span}}+\lambda_{\text{2}}\mathcal{L}_{\text{KL}}+\lambda_{\text{3}}\mathcal{L}_{\text{seg}},$ where $\mathcal{L}_{\text{LM}}$ is the standard language-generation loss. The BIO head precisely extracts segmentation spans, while the Focal-Tversky loss, biased toward recall ($\alpha\!=\!0.7$) and precision ($\beta\!=\!0.3$), encourages sharper, high-IoU masks at fine-grained image regions.
@@ 5 Experiments
We use a pre-trained LMM, LLaVA-7B, and LLaVA-llama2-13B (liu2023llava, ) as a backbone for PLUM (Sec. 4). PLUM follows the consecutive two-stage finetuning process: (1) PLUM is first finetuned with a randomly sampled mixture of PACO-LVIS (ramanathan2023paco, ), Pascal Parts (chen2014pascalpart, ), PartImageNet (he2022partimagenet, ), COCO-Stuff (caesar2018cocostuff, ), ADE20k (zhou2017scene, ), RefCOCO line of datasets (kazemzadeh-etal-2014-referitgame, ), and a VQA dataset from LLaVA (llava_instruct_150k) and a Reasoning Segmentation (lai2024lisa, ) training dataset. This setting is similar to the previous line of segmentation-enabled LMMs (lai2024lisa, ; rasheed2024glamm, ; ren2024pixellm, ), and we refer to the stage-1 checkpoint of PLUM as the zero-shot (or pretrained) baseline throughout this paper. (2) To further finetune PLUM on our Partonomy training dataset, we take Partonomy-PACO, -PartImageNet and -PascalParts to construct a training split and a validation split. Note, we do not use the Partonomy-Core split as training data and use it only as evaluation data. We refer to the Appendix for additional details on the hyperparameter settings and training details.

### v3

@@ Abstract
Real-world objects are composed of distinctive, object-specific parts. Identifying these parts is key to performing fine-grained, compositional reasoning—yet, large multimodal models (LMMs) struggle to perform this seemingly straightforward task. In this work, we introduce Partonomy, an LMM benchmark designed for pixel-level part grounding. We construct Partonomy from existing part datasets and our own rigorously annotated set of images, encompassing 862 part labels and 534 object labels for evaluation. Unlike existing datasets that simply ask models to identify generic parts, Partonomy uses specialized concepts (e.g., agricultural airplane), and challenges models to compare objects’ parts, consider part-whole relationships, and justify textual predictions with visual segmentations. Our experiments demonstrate significant limitations in state-of-the-art LMMs (e.g., LISA-13B achieves only 5.9% gIoU), highlighting a critical gap in their part grounding abilities. We note that existing segmentation-enabled LMMs (segmenting LMMs) have two key architectural shortcomings: they use special [SEG] tokens not seen during pretraining which induce distribution shift, and they discard predicted segmentations instead of using past predictions to guide future ones. To address these deficiencies, we propose Plum, a novel segmenting LMM that uses span tagging instead of segmentation tokens and that conditions on prior predictions in a feedback loop. We find that pretrained Plum outperforms existing segmenting LMMs on reasoning segmentation, VQA, and visual hallucination benchmarks. In addition, Plum finetuned on our proposed Explanatory Part Segmentation task is competitive with segmenting LMMs trained on significantly more segmentation data. Our work opens up new avenues towards enabling fine-grained, grounded visual understanding in LMMs. The code and data are publicly available at: https://github.com/AnselBlume/partonomy
@@ 1 Introduction
Figure 1: The Partonomy dataset evaluates LMMs’ part understanding through the Explanatory Part Segmentation task. Given an input image, a segmentation-enabled LMM selects a textual explanation and generates part segmentation masks which serve as textual and visual rationale for its answer choice. Our question-answer mutation framework generates challenging answer choices by predicting part co-occurrence and by selecting parts from confusable objects.
Real-world objects can be decomposed into distinctive parts. A banana boat (Fig.˜2), for instance, consists of a seating tube, a handle, a hull, and an inflation valve. Such parts characterize each concept, differentiating one object from another. The ability to recognize and distinguish between parts is an important element of holistic object understanding, with applications ranging from explainable object recognition blume2024miracle ; kim-ji-2024-finer ; liu2023llava ; dai2023instructblip ; zhu2023minigpt , to part-based novel concept design design2025 , and robotic manipulation xu2023creativerobottooluse ; liu2024composable ; jiang2024robot . Decomposing objects into their key building blocks allows models to reason about objects at a granular level kim-ji-2024-finer ; liu2024democratizing , allowing for more complex and nuanced interactions.
@@
To quantify the part recognition abilities of LMMs, we propose Explanatory Part Segmentation, a task that assesses LMMs’ ability to recognize object parts, associate objects with their distinctive parts, and use these grounded parts to predict object labels. We then introduce Partonomy, a comprehensive benchmark for the Explanatory Part Segmentation task. We construct Partonomy from existing part segmentation datasets he2022partimagenet ; chen2014pascalpart ; ramanathan2023paco and our manually-annotated evaluation dataset of 1K specialized object-centric images with complex part annotations. This subset, Partonomy-Core, contains 862 distinct part labels—more than any existing part datasets (Table˜1).
We then note two shortcomings of existing segmenting LMMs’ architectures lai2024lisa ; rasheed2024glamm ; ren2024pixellm . First, they always rely on special [SEG] tokens not seen during pretraining, potentially hindering downstream performance by introducing distribution shift. Second, these segmenting LMMs discard their predictions after each output, missing the opportunity to incorporate prior information contained by the masks they predicted during the decoding process. This design is in contrast to modern generative frameworks, which condition future predictions on past ones vaswani2017attention ; sohl2015thermodynamics . Based on these observations, we propose Plum, a Part-Level Understanding LMM. Plum uses a text span tagging module to avoid special segmentation tokens that induce distribution shift from the pre-trained LLM, and employs a mask feedback mechanism to condition on past predictions (Section˜4). Our results show that pretrained Plum retains its general reasoning abilities far better than other segmenting LMMs, achieving stronger zero-shot segmentation performance and competitive finetuned performance to models trained on significantly more segmentation data.
@@ 2 Related Work
Figure 2: An example of PLUM’s part understanding compared to recent segmenting LMMs trained on part data.
LMMs such as LISA (lai2024lisa, ) and GLaMM (rasheed2024glamm, ) have demonstrated the ability to generate text and grounded segmentation masks. Despite being trained on part-level segmentation datasets such as PACO (ramanathan2023paco, ) and Pascal-Part (pascal-voc-2007, ), they struggle to exhibit a part-level understanding of visual concepts. While they demonstrate the ability to understand complex textual instructions (lai2024lisa, ; rasheed2024glamm, ; ren2024pixellm, ; yuan2024osprey, ), current LMMs fail to relate concept-indicative parts to their wholes, as shown in Fig.˜2. Frequently, they fail to generate the specialized segmentation token (e.g., [SEG]) added to their vocabulary, leading to no masks being generated for the parts. Even state-of-the-art LMMs that seemingly “reason” struggle to establish attributive relationships between objects and parts, implying that current models and datasets lack the coverage and capacity to handle complex part understanding and grounding tasks. This observation motivates the proposal of our new Explanatory Part Segmentation task, which requires LMMs to segment objects’ parts (i.e. producing visual rationale) while generating the corresponding text rationale.
@@ 3.1 Task Overview
To evaluate these elements of part understanding, we define the Explanatory Part Segmentation task. In this task, a model is provided with an image and a question about an object’s parts (e.g. “What visible parts does the agricultural plane in the image have?” Fig.˜1). The model must then select the best response and generate segmentation masks for the corresponding parts to explain its selection (e.g. “The agricultural plane has wings, a propulsion component, and a spraying rig”). Motivated by our characterization of part understanding, we define three classes of questions (Fig.˜1):
@@
Datasets # Object Labels # Part Labels # Object-Part Labels # Images # Seg. Masks
Object‐Centric
Images
Object
Domain
PASCAL‐Part 20 30 193 10,103 111,960 C, O
PartImageNet 158 14 14 24,000 112,000 C, O
PartImageNet++ 1000 818 3,308 100K 406.4K C, O
PACO 75 200 456 84,027 641,000 C
Partonomy ‐Core 534 862 1,976 1,068 4,968 C, O, S
@@ 3.3 Evaluation
We evaluate answer selection via accuracy (random == 20% for 5 answer choices) for both part prediction (all questions) and object prediction (Part-Whole Reasoning questions). However, some wrong answer choices are better than others. Precision and recall capture the similarity of the predicted parts to the ground truth set, and we adopt these as more fine-grained measurements of part recognition by the LMMs.
@@ 4 Plum: Part-Level Understanding LMM
We find that existing LMMs are unable to accurately identify parts in an image. Even segmenting LMMs (lai2024lisa, ; rasheed2024glamm, ; ren2024pixellm, ) trained on part segmentation datasets such as PACO (ramanathan2023paco, ) and Pascal-Part (chen2014pascalpart, ) exhibit poor performance on part-level segmentation (Figs.˜4 and 2). We identify two key architectural deficiencies of segmenting LMMs: (1) They rely on special tokens for segmentation (e.g. [SEG] or <p></p>). These tokens are not seen during pretraining, so we hypothesize that their addition to the vocabulary and subsequent finetuning perturbs models’ original token distributions (Table˜5). (2) They discard prior mask predictions when segmenting in sequence, conditioning only on past text during generation. Incorporating prior mask predictions would likely help maintain consistency and better localize future predictions (Fig.˜5(a)).
@@
Let $h_{i}^{L}\!\in\!\mathbb{R}^{d}$ be the VLM’s last-layer embedding of token $i$ $(i=1,\dots,N)$ of the output sequence. We process these embeddings along two complementary pathways: (i) the Span Extractor, a bidirectional self-attention block that tags beginning (B), inside (I), and outside (O) (ramshaw1999text, ) positions of tokens to segment, and (ii) a projection head that maps B/I embeddings into “mask queries,” regularized by KL divergence. An overview is given in Fig.˜3.
@@
During inference, contiguous $\textit{B}\!\to\!\textit{I}$ chains are greedily merged to form text spans that are projected into segmentation queries. Note, we also enable users of Plum to override the automatic tags with manual span selection, enabling interactive, interpretable “highlight-to-segment” behavior.
@@
Let $\mathcal{S}=\{(i_{s},j_{s})\}_{s=1}^{N_{+}}$ be the set of contiguous $\textit{B}\to\textit{I}$ spans produced by the span extractor, where $i_{s}$ and $j_{s}$ are the start and end token indices of span $s$, and $N_{+}=|\mathcal{S}|$ is the total number of such spans in the sequence. For every span token $k\in[i_{s},j_{s}]$ we obtain a “mask-query” vector $q_{k}=g(h_{k}^{L})\in\mathbb{R}^{m}$, with $g(\cdot)$ a learned MLP projection. To keep the span representations close to the pre-trained backbone VLM’s manifold, we pool the last-layer embeddings of each span [^2] We use mean pooling; any differentiable pooling operator is admissible. and impose a Gaussian KL penalty against the corresponding frozen teacher embedding $t^{L}_{i_{s}:j_{s}}$: $\small\mathcal{L}_{\text{KL}}=\frac{1}{N_{+}}\sum_{s=1}^{N_{+}}\frac{\bigl\|h^{L}_{i_{s}:j_{s}}-t^{L}_{i_{s}:j_{s}}\bigr\|_{2}^{2}}{2\sigma^{2}}.$ This term is applied only to B/I spans, preventing their hidden states from drifting away from the original language-representation space and thereby preserving the VLM’s textual reasoning ability.
@@
To incorporate previously predicted masks into the mask decoding process, we inject feature-wise linear modulation (FiLM) (perez2018film, ) layers into the SAM decoder’s mask encoder (Fig.˜3). These layers allow us to encode the mask while conditioning on prior text spans, providing semantics beyond a raw binary mask. We use this modified mask encoder to encode each prior mask into a stack of text-enhanced feature maps. The stack of feature maps is pooled into a single feature map via patch-wise attention pooling (over the stack dimension), with a learned feature map providing an attentional query for each patch. This pooled feature map representing all prior predicted masks is fed into the mask decoder (along with the pooled text embeddings) to generate the next mask.
@@
Tagged token embeddings $q_{k}$ are average pooled and passed to the mask decoder, generating a mask $\hat{M}_{i}$. With ground-truth mask $M_{i}$ we adopt the Focal-Tversky loss (abraham2019novel, ), $\mathcal{L}_{\text{seg}}=\frac{1}{N_{+}}\!\sum_{y_{i}\neq\textsc{O}}\mathcal{L}_{\text{FT}}(M_{i},\hat{M}_{i})$. Focal-Tversky loss is a generalized version of the DICE loss (sudre2017generalised, ). This gives the overall objective equal to $\mathcal{L}=\mathcal{L}_{\text{LM}}+\lambda_{\text{1}}\mathcal{L}_{\text{span}}+\lambda_{\text{2}}\mathcal{L}_{\text{KL}}+\lambda_{\text{3}}\mathcal{L}_{\text{seg}}+\lambda_{\text{4}}\mathcal{L}_{\text{BCE}},$ where $\mathcal{L}_{\text{LM}}$ is the standard language-generation loss and $\mathcal{L}_{BCE}$ is per-pixel binary cross-entropy as adopted from (lai2024lisa, ). The BIO head precisely extracts segmentation spans, while the Focal-Tversky loss, biased toward recall ($\alpha\!=\!0.7$) and precision ($\beta\!=\!0.3$), encourages sharper, high-IoU masks at fine-grained image regions. For additional details on the hyperparameter setting, refer to §A.1
@@ 5 Experiments
We use a pre-trained LMM, LLaVA-7B, and LLaVA-llama2-13B (liu2023llava, ) as backbones for PLUM (Sec. 4). PLUM follows the consecutive two-stage finetuning process: (1) PLUM is first finetuned with a randomly sampled mixture of PACO-LVIS (ramanathan2023paco, ), Pascal Parts (chen2014pascalpart, ), PartImageNet (he2022partimagenet, ), COCO-Stuff (caesar2018cocostuff, ), ADE20k (zhou2017scene, ), the RefCOCO line of datasets (kazemzadeh-etal-2014-referitgame, ), a VQA dataset from LLaVA (llava_instruct_150k), and a Reasoning Segmentation (lai2024lisa, ) dataset. This setting is similar to the previous line of segmentation-enabled LMMs (lai2024lisa, ; rasheed2024glamm, ; ren2024pixellm, ), and we refer to the stage-1 checkpoint of PLUM as the zero-shot (or pretrained) baseline throughout this paper. (2) To further finetune PLUM on our Partonomy training dataset, we take Partonomy-PACO, -PartImageNet and -PascalParts to construct a training split and a validation split. Note, we do not use the Partonomy-Core split as training data and use it only as evaluation data. We refer to the Appendix for additional details on the hyperparameter settings and training details.
Baselines
To evaluate PLUM’s proposed changes, we use LISA lai2024lisa , GLaMM rasheed2024glamm , and PixelLM ren2024pixellm as our primary baselines. All of these models use LLaVA as the base LMM and use SAM-style decoders to generate segmentation masks liu2023llava ; kirillov2023segment .
For segmentation, we also evaluate X-Decoder, SEEM, and Grounded SAM 2 as general open-vocabulary segmentation models zou2023generalized ; zou2023segment ; ren2024grounded ; ravi2024sam2 . As they do not understand question-based prompts, we provide them with the ground truth (gt) parts to segment individually. We also include SegLLM, a segmenting LMM based on LLaVAv1.5 and HIPIE liu2024llavav1.5 ; wang2023hipie . HIPIE is a decoder built for multi-scale and part segmentation, and SegLLM is trained to perform multi-round segmentation on parts, allowing it to refer back to previously predicted masks. The similarity of this mechanism to our feedback loop motivates us to include SegLLM as a baseline, despite the imperfect comparison due to its use of a newer LMM and different mask decoder.
For text evaluations, we include a random baseline (which randomly selects answers) to situate the models’ part precision and recall. GPT-4o [^3] Specifically, gpt-4o-2024-08-06., a frontier model, provides an upper bound on performance. GPT-4o has an advantage as it must be provided with all four answer choices at once, allowing it to take advantage of shortcuts not available to the other models (like identifying the base answer from which the wrong answer choices are generated).